FTMW SPECTROSCOPY OF AlSH AND AlSD

\underline{MASARU FUKUSHIMA}^{a}^{b} and MICHAEL C. L. GERRY, Department of Chemistry, The University of British Columbia, 2036 Main Mall, Vancouver, BC, Canada, V6T 1Z1.

We have observed AlSH and AlSD for the first time using the technique of Fourier transform microwave spectroscopy. AlSH (AlSD) was produced in supersonic free jet expansions by reaction of H_{2}S (D_{2}S) with laser ablated Al atoms. The spectral search was guided by an ab initio computional prediction performed prior to the experiments. The lowest frequency rotational transition, 1_{01} - 0_{00}, was measured and the hyperfine structure due to the nuclear spin of ^{27}Al was analyzed. By comparing the obtained rotational constant, \frac{1}{2} (B+C), with the computational prediction, it was concluded that AlSH has a bent structure with a bond angle close to 90^{∘}. Using the nuclear electric quadrupole and nuclear spin-rotation coupling constants obtained, the electronic structure of AlSH was investigated.

aPermanent address : Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

bSupport by Hiroshima City University President-Designated Grant for Long-Term Research Abroad for Young Researchers is gratefully acknowledged.